OpenStudy (anonymous):

Help If f(x)=1/2x+14, then f^-1(x)=

Nnesha (nnesha):

f(x) is same as y so first replace f(x) with y \[\large\rm \color{ReD}{y}=\frac{ 1 }{ 2\color{blue}{x} }+14\]then switch x and y \[\large\rm \color{blue}{x}=\frac{ 1 }{ 2\color{reD}{y}}+14\] now that equation for y

Nnesha (nnesha):

btw it's \[\frac{ 1 }{ 2x }+14 ~~~~~or~~~~~~~~~~\frac{ 1 }{ 2x+14 }\]

Nnesha (nnesha):

???